OpenStudy (anonymous):

A loan of Rs. 2000 is to be paid back in three equal annual installments. How much is each installment approximately, if the interest is compounded annually at 12 % p.a.?

OpenStudy (anonymous):

@shrutipande9

OpenStudy (shrutipande9):

@nurali can definitely help u well

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

@ganeshie8

ganeshie8 (ganeshie8):

are you ready for a short derivation ? :)

OpenStudy (anonymous):

hmmmm.....

ganeshie8 (ganeshie8):

derivation is easy, but may be lets just memorize the EMI formula as you don't have much time for exam : http://emicalculator.net/wp-content/uploads/2011/06/emiformula.png?fe8349

OpenStudy (anonymous):

k

ganeshie8 (ganeshie8):

\[\large E =P.r . \dfrac{(1+r)^n}{(1+r)^n-1} \] E =periodic payment P = loan amount = 2000 r = periodic interest rate = 12% = 0.12 n = number of periods = 3 plug and chug

ganeshie8 (ganeshie8):

seen that formula before ?

OpenStudy (anonymous):

26000

ganeshie8 (ganeshie8):

ugh, how can u pay 26,000 every year when for a loan of just 2000 ?

OpenStudy (anonymous):

i might have but forgetten.i will always search for shortcuts

ganeshie8 (ganeshie8):

\[\large E =2000(0.12) \dfrac{(1+0.12)^3}{(1+.12)^3-1}\]

ganeshie8 (ganeshie8):

simplify ^^

OpenStudy (anonymous):

oh! n=1 i used

ganeshie8 (ganeshie8):

I see... you're paying back in 3 annual installments, not 1

OpenStudy (anonymous):

i got 840

ganeshie8 (ganeshie8):

looks you're rounding the numbers too much

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

832.7